The length of a side of an equilateral triangle is the same as the length of a rectangle and the width of the rectangle is 2 inches less than its length. if the primeter of the triangle is 4 inches less than the primeter of the rectangle, what are the dimensions of the rectangle?
Answer by ankor@dixie-net.com(22740)

The length of a side of an equilateral triangle is the same as the length of a rectangle and the width of the rectangle is 2 inches less than its length.
if the perimeter of the triangle is 4 inches less than the perimeter of the rectangle, what are the dimensions of the rectangle?
:
Let x = the length of the side of the triangle and the length of the rectangle
then
(x-2) = the width of the rectangle
:
"the perimeter of the triangle is 4 inches less than the perimeter of the rectangle,"
3x = 2x + 2(x-2) - 4
3x = 2x + 2x - 4 - 4
3x = 4x - 8
3x - 4x = -8
-x = -8
therefore
x = 8 inches is the length of the rectangle (and the side of the triangle)
then
8 - 2 = 6 in is the width of rectangle
:
8 by 6 is the dimensions of the rectangle
:
:
See if that checks out:
Triangle 3(8) = 24 in
Rect: 2(8)+2(6) = 28 in
-----------------------
difference: 4 inches